one kg of air expand at a constant temperature from a pressure of 820 kpa and a volume of 3 m³ to a pressure of 220 kpa. determine.




A. Change of Internal Energy




B. Change of Enthalpy

1
Expert's answer
2022-02-20T15:48:56-0500

"W=P\u2206V"

"W=P(V_2-V_1)"

"\frac{P_1}{P_2}=\frac{V_1}{V_2}\\V_2=\frac{P_2}{P_1}V_1"

"V_2=\frac{820}{220}\times3=11.18m^3"


"W=P(V_2-V_1)=820\times10^3\times(11.18-3)=6.7076KJ"

∆Q=mRT

"∆Q=1\times8.31\times273=2.278KJ"

Thermodynamic first law

"\u2206U=\u2206Q-\u2206W"


"\u2206U=(2.278-6.7076)KJ=-4.4296KJ"

Enthalpy change

"\u2206H=U+P\u2206V"


"\u2206H=-4.4296+6.7076=2.278KJ"